# A Minkowski space embedding to understand Markov models dynamics

### v1

@@ V Summary and outlook
The embedding coordinates structure the space of Markov models in terms of equilibrium and nonequilibrium (i.e. symmetric and antisymmetric) components. The equilibrium coordinates are expressed in terms of the system’s activities and kinetic forces while the nonequilibrium coordinates coincide with the dynamical equivalence classes and are expressed in terms of the system’s currents and affinities. The symmetrized Kullback-Leibler divergence and resulting coordinates measure the time and spatial asymmetry between models.
This framework provides a systematic way to analyze the space of Markov models and to gain insights into the behavior and design of mesocopic devices. For example, the Minkowski embedding allows to explore a system’s response in the directions ’orthogonal’ to the dynamical equivalence classes, where additional symmetries are expected to appear. The embedding can be used to determine the minimum parameters needed to describe model data or to identify the set of governing parameters for model prediction and coarse graining [4, 28]. The space-time structure could also help untangle other important phenomena such as dynamical or nonequilibrium phase transitions.
Disclaimer. This paper is not intended for journal publication.
Appendix: Derivation of the embedding coordinates
The first step to obtain Eq. (4) is to calculate the log ratios $\ln P_{ij}/G_{ij}$ that appear in Eq. (3). Using the thermodynamic decomposition (1) with $P=s[E\circ Z]$ and $G=s[E^{\prime}\circ Z^{\prime}]$, the log ratios take the form
$\ln\frac{P_{ij}}{G_{ij}}=\ln\frac{\rho^{\prime}}{\rho}+\ln\left(\frac{\alpha_{j}^{\prime}}{\alpha_{i}^{\prime}}\frac{\alpha_{i}}{\alpha_{j}}\right)+\ln\frac{Z_{ij}}{Z^{\prime}_{ij}}+\ln\frac{E_{ij}}{E^{\prime}_{ij}}\,.$ (13)
The first term $\ln\rho^{\prime}/\rho$ vanishes when inserted into expression (3) since $\sum\pi_{i}P_{ij}=\sum\nu_{i}G_{ij}=1$. The second term vanishes as well when averaged over a stochastic dynamics (see for example Lemma 4.3 (iii) in ref. [16]). The third term in Eq. (13) reads
@@
To calculate the last term $\ln E_{ij}/E^{\prime}_{ij}$, remember that $E_{ij}=(1/\lambda)(\gamma_{i}/\gamma_{j})\sqrt{P_{ij}P_{ji}}$ where $\lambda$ and $\gamma$ are the Perron eigenvalue and eigenvector of $\sqrt{P_{ij}P_{ji}}$, and similarly for $E^{\prime}$. Here also the terms $\ln\lambda^{\prime}/\lambda$ and $\ln(\gamma^{\prime}_{j}\gamma_{i}/\gamma^{\prime}_{i}\gamma_{j})$ will vanish when averaged over $\sum\pi_{i}P_{ij}-\nu_{i}G_{ij}$. The remaining factor then becomes

### v3

@@ V Summary and outlook
The embedding coordinates structure the space of Markov models in terms of equilibrium and nonequilibrium (i.e. symmetric and antisymmetric) components. The equilibrium coordinates are expressed in terms of the system’s activities and kinetic forces while the nonequilibrium coordinates coincide with the dynamical equivalence classes and are expressed in terms of the system’s currents and affinities. The symmetrized Kullback-Leibler divergence thus measures the difference in temporal and spatial structures between models.
This framework provides a systematic way to analyze the space of Markov models and to gain insights into the behavior and design of mesocopic devices. For example, the Minkowski embedding allows to explore a system’s response in the directions ’orthogonal’ to the dynamical equivalence classes, where additional symmetries are expected to appear. The embedding can be used to determine the minimum parameters needed to describe model data or to identify the governing parameters for model prediction and coarse graining [4, 29]. The space-time structure could also help untangle important phenomena such as dynamical or nonequilibrium phase transitions.
Appendix A: Derivation of the embedding coordinates
The first step to obtain Eq. (4) is to calculate the log ratios $\ln P_{ij}/Q_{ij}$ that appear in Eq. (3). For notational simplicity, we use $P^{e}=E$ in this section. Using the thermodynamic decomposition (1) with $P=s[E\circ Z]$ and $Q=s[E\circ Z^{\prime}]$, the log ratios take the form
$\ln\frac{P_{ij}}{Q_{ij}}=\ln\frac{\rho^{\prime}}{\rho}+\ln\left(\frac{\alpha_{j}^{\prime}}{\alpha_{i}^{\prime}}\frac{\alpha_{i}}{\alpha_{j}}\right)+\ln\frac{Z_{ij}}{Z^{\prime}_{ij}}+\ln\frac{E_{ij}}{E^{\prime}_{ij}}\,.$ (12)
The first term $\ln\rho^{\prime}/\rho$ vanishes when inserted into expression (3) since $\sum\pi_{i}P_{ij}=\sum\nu_{i}Q_{ij}=1$. The second term vanishes as well when averaged over a stochastic dynamics (see for example Lemma 4.3 (iii) in ref. [18]). The third term in Eq. (12) reads
@@
To calculate the last term $\ln E_{ij}/E^{\prime}_{ij}$, remember that $E_{ij}=(1/\lambda)(\gamma_{i}/\gamma_{j})\sqrt{P_{ij}P_{ji}}$ where $\lambda$ and $\gamma$ are the Perron eigenvalue and eigenvector of $\sqrt{P_{ij}P_{ji}}$, and similarly for $E^{\prime}$. Here also the terms $\ln\lambda^{\prime}/\lambda$ and $\ln(\gamma^{\prime}_{j}\gamma_{i}/\gamma^{\prime}_{i}\gamma_{j})$ will vanish when averaged over $\sum\pi_{i}P_{ij}-\nu_{i}Q_{ij}$. The remaining factor then becomes